OpenStudy (unklerhaukus):

solve the differential equation \[\frac{\text dy}{\text dx}-y=xy^{1/2}\]

OpenStudy (unklerhaukus):

\[\frac{\text dy}{\text dx}-y=xy^{1/2}\] \[\text{let }v=y^{1-1/2}=y^{1/2}\]\[\text dv=\frac 1{2y^{1/2}}\text dy\]\[\text dy={2v}\text dv\] \[2v\frac{\text dv}{\text dx}-v^2=xv\]\[\frac{\text dv}{\text dx}-\frac 12 v=\frac x2\] \[R(x)=e^{\int -\frac 12\text dx}=e^{-\frac x2}\] \[\frac{\text d}{\text dx}\left(ve^{-\frac x2}\right)=\frac{xe^{-\frac x2}}2\]\[ve^{-\frac x2}=\int\frac{xe^{-\frac x2}}2\text dx\]

OpenStudy (anonymous):

sqrt(y(x))+2+x-exp((1/2)*x)*_C1 = 0

OpenStudy (unklerhaukus):

how did you ?

OpenStudy (unklerhaukus):

@timo86m did you use any of my working ? how did you get that result/

OpenStudy (unklerhaukus):

\[ve^{-\frac12}=\frac 12\int xe^{-\frac x2}\text dx\]

OpenStudy (ash2326):

It's correct:)

OpenStudy (unklerhaukus):

i guess i have to integrate by parts now

OpenStudy (ash2326):

Yeah:) use by parts

OpenStudy (unklerhaukus):

\(\int uv'=\left.uv\right|-\int vu'\)

OpenStudy (ash2326):

Yeah:D you're doing great

OpenStudy (unklerhaukus):

\[u=x\qquad u'=1\]\[v'=e^{-x/2}\qquad v=\frac{-e^{-x/2}}{2}\]

OpenStudy (unklerhaukus):

\[ve^{-\frac x2}=\left.-\frac{xe^{-x/2}}{2}\right|+\int \frac{e^{-x/2}}2\text dx\]\[ve^{-\frac x2}=\left.\frac{-xe^{-x/2}}{2}\right|-{e^{-x/2}}+c\]

OpenStudy (lalaly):

correction .. \[\large{}v' =e^{-\frac{x}{2}} \] \[\large{v=-\frac{e^{-\frac{x}{2}}}{\frac{1}{2}}}=-2e^{-\frac{x}{2}}\]

OpenStudy (unklerhaukus):

Ah!

OpenStudy (unklerhaukus):

Thank YOU

OpenStudy (lalaly):

:D

OpenStudy (unklerhaukus):

\[ve^{-\frac x2}=\left.-{2}{xe^{-x/2}}\right|+2\int {e^{-x/2}}\text dx\] \[ve^{-\frac x2}=\left.-2xe^{-x/2}\right|-4{e^{-x/2}}+c\]

OpenStudy (lalaly):

yes and v=sqrt(y)

OpenStudy (unklerhaukus):

what to i do with the limit \(\left. \right|\)

OpenStudy (unklerhaukus):

\[\sqrt y={-2x}-4+ce^{x/2}\]

OpenStudy (unklerhaukus):

This is the answer in the back of the book _______________________\[\sqrt y=-2-x+ce^{x/2}\]_______________________

OpenStudy (lalaly):

can u give me a min i want to try

OpenStudy (lalaly):

v=sqrt(y) dv=1/(2sqrt(y)) \[v'-\frac{v}{2}=\frac{x}{2}\] \[\large{I=e^{\int\limits{-\frac{1}{2}}dx}}=e^{-\frac{x}{2}}\]multiplying the DE by I \[e^{-\frac{x}{2}}v'-\frac{e^{-\frac{x}{2}}}{2}v=\frac{xe^{-\frac{x}{2}}}{2}\]\[ve^{-\frac{x}{2}}=\frac{1}{2} \int\limits{xe^{-\frac{x}{2}}dx}\]now integrating u=x dv=e^(-x/2)dx du=dx v=-2e^(-x/2) \[ve^{-\frac{x}{2}}=\frac{1}{2}(-2xe^{-\frac{x}{2}}+2\int\limits\limits\limits{e^{-\frac{x}{2}}dx})\]\[ve^{-\frac{x}{2}}=\frac{1}{2}(-2xe^{-\frac{x}{2}}-4e^{-\frac{x}{2}}+C)\]now divide everything by e^(-x/2) \[v=\frac{1}{2}(-2x-4+ce^{\frac{x}{2}})\]\[v=-x-2+\frac{c}{2}e^{\frac{x}{2}}\]

OpenStudy (lalaly):

c/2 is any constant so it can be just C

OpenStudy (lalaly):

no need for any limits in integration ,,, and u forgot the 1/2 u put at the beginning

OpenStudy (lalaly):

@UnkleRhaukus

OpenStudy (lalaly):

so the answer is \[\sqrt{y}=-x-2+ce^{\frac{x}{2}}\]

OpenStudy (unklerhaukus):

yeah forgot about the \(\frac 12\) when i integrated by parts thanks for all your work

OpenStudy (lalaly):

anytime

OpenStudy (unklerhaukus):

in the integration by parts bit we are ment to drop the limit because we are adding the arbitrary constant of integration/ is this right? \(\int\limits_a^b uv'=\left.uv\right|_a^b-\int\limits_a^b vu'\) \(\int uv'=uv-\int\limits_a^b vu'+c\)

OpenStudy (unklerhaukus):

?

OpenStudy (lalaly):

yes its an indefinite integral